OpenStudy (anonymous):

factor this polynomial completely x^2-12x-36

OpenStudy (anonymous):

hold on one minute

OpenStudy (anonymous):

1.1 Factoring x2-12x+36 The first term is, x2 its coefficient is 1 . The middle term is, -12x its coefficient is -12 . The last term, "the constant", is +36 Step-1 : Multiply the coefficient of the first term by the constant 1 • 36 = 36 Step-2 : Find two factors of 36 whose sum equals the coefficient of the middle term, which is -12 . -36 + -1 = -37 -18 + -2 = -20 -12 + -3 = -15 -9 + -4 = -13 -6 + -6 = -12 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, -6 and -6 x2 - 6x - 6x - 36 Step-4 : Add up the first 2 terms, pulling out like factors : x • (x-6) Add up the last 2 terms, pulling out common factors : 6 • (x-6) Now add up the four terms of step 3 : (x-6) • (x-6) Which is the desired factorization or (x - 6)2 = 0 Step 2 : Solve (x-6)2 = 0

OpenStudy (anonymous):

trayall, shouldn't it rather be \[x^2 -12x + 36\]?